OpenStudy (anonymous):

negative 1 over 2 x + 2 = −x + 7

OpenStudy (anonymous):

\[\frac{-1}{2x}+2=-x +7\] To solve for x, you need to isolate the variables. Try to combine the constants (the numbers) first.

OpenStudy (anonymous):

\[\frac{-1}{2x+2}=-x+7\] \[-1=(2x+2)(-x+7)\] solve the quadratic